Murphy Inc. operates in the automotive parts business. Murphy is considering expanding. Currently, Murphy has a before tax cost of debt of 6%, and an effective tax rate of 40%. Murphy's capital structure consists of 40% debt and 60% equity. Murphy has identified the following pure play company: Brown Inc. who operates in a similar automotive parts line of business. Brown Inc. has a before tax cost of debt of 5%, a debt to equity ratio of 0.50, and an effective tax rate of 30%. Its stock has as much systematic risk as the market portfolio. The market risk premium is 4%, and the risk free rate of interest is 3%.

- Determine the appropriate cost of capital (WACC) for the Murphy Inc should use for evaluating an expansion in their automotive parts division. Assume the expansion project will have a debt to equity ratio similar to that of Murphy. Hint: Unlever Brown's beta, Re-lever the beta using Murphy's's debt to equity ratio and tax rate. Then determine the cost of equity for the expansion project. Insert the cost of equity in the WACC and determine the WACC for the expansion project
Answer #1

Cost of Equity = Rf + beta * (Risk Premium)

= 3% + 1 * (4%)

= 7%

Beta of Market is always one. So it is given the beta is same as market. we have taken beta as 1.

Pre - tax Cost of Debt = 5%

Debt Equity = 0.50 / 1

Total Weight = 0.50 + 1 = 1.5

Weight of Equity = 1 / 1.5 = 0.67

Weight of Debt = 0.50 / 1.5 = 0.33

Cost of Capital = (Cost of Equity * Weight of Equity) + (Cost of Debt after tax * Weight of Debt)

= 7% * 0.67 + 5% * (1-0.30) * 0.33

= 5.85%

WACC of 5.85% can be used.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.